# Cancellation of phonon hopping in trapped ions by modulation of the trap potential

### v1

@@ Abstract
The local modes of trapped ions can be used to construct an analog quantum simulator and a digital quantum computer. However, the control of the phonon hopping remains difficult because it proceeds among all the local modes through the Coulomb coupling. We propose a method to cancel the phonon hopping among a given set of local modes by applying a sequence of phase shift gates implemented through the modulation of the trap potential. We analyze the error scaling in the algorithm to treat three or more modes and show that the error can be suppressed by repeating the pulse sequence. The duration of the phase shift gate in the present method can be as short as a few microseconds, which is an order of magnitude faster than the laser-based method. This short duration of the phase shift gate facilitates the suppression of the gate error. The present method can be applied to the simulation of bosonic systems as well as to the continuous variable encoding of quantum computing using trapped ions.
pacs:
@@ I INTRODUCTION
The local mode scheme can be used to realize the quantum simulation of, e.g., a Bose–Hubbard model [3] and a Jaynes–Cummings–Hubbard model [4], and can also be used to implement a quantum walk [6] and a boson sampling [7]. Moreover, it has been shown that a logical qubit in the CV encoding called the Gottesman–Kitaev–Preskill (GKP) state can be constructed using a local mode [8, 9, 10].
However, in the local mode scheme, the manipulation of entanglement among two or more local modes remains difficult because the phonon hopping, a source of entanglement, originating from the Coulomb interaction cannot be turned off. On the other hand, in the collective mode scheme [11], the modes can be entangled using a beam-splitter gate, which can be switched on only by irradiating the ions with laser [11] or by applying the spatially and temporally varying electric potential [12].
There are several methods for the control of the phonon hopping. In the explanation below, we refer to a system having two local modes for simplicity, but the discussion can be extended to a system having more than two local modes. The first method is called the phonon blockade [13], where two local modes become off-resonant by the irradiation of one of the two local modes with laser light so that the phonon hopping can be suppressed. The method protects the phonon state of one mode, but disturb that of the other mode being excited by the laser light.
@@
In the present study, we adopt the third method and propose to implement the $\pi$-phase shift by modulating the trap potential using an electric pulse applied to the trap electrodes and show that the fidelity of the DD can be improved. The duration of $P$ can be a few microseconds, which is shorter by an order of magnitude than that can be achieved by the resonant sideband pulse approach. It should be noted that the present method can be applied to a system having any number of phonons. We also show that the original DD approach  [7] needs to be modified when the number of modes becomes larger than two even when the duration of $P$ is negligibly small.
@@ II.1 A sequence of phase shifts for the dynamical decoupling
We consider a chain of ions and assume that the frequencies of the oscillation along the radial direction $\tilde{x}$, perpendicular to the trap axis, are a few MHz while the hopping rate is a few kHz. This condition allows us to adopt the local-mode picture along $\tilde{x}$ axis. Therefore, if the number of trapped ions is $M$, we consider $M$ local modes and denote the set of those modes as $\mathcal{U}=\{q_{j};0\leq j\leq M-1\}$. The scaled coordinate of the $j$-th mode $x_{j}$ is represented using the annihilation and creation operators as $x_{j}=(a_{j}^{\dagger}+a_{j})/\sqrt{2}$. As shown in Appendix. A, when the distance between adjacent ions is much larger than the displacement of the respective ions from their equilibrium positions, the Hamiltonian is composed of the on-site interaction $H_{0}$ and the Coulomb interaction $H_{C}$ as
@@
where $\omega_{0}$ in Eq. (2) is the secular frequency obtained under the influance of the Coulomb interactions. The rotating-wave approximation (RWA) has been applied to $H_{C}$ so that the fast oscillating terms such as $a_{j}a_{k}+a_{j}^{\dagger}a_{k}^{\dagger}$ can be omitted. The Hamiltonian in the interaction picture with respect to $H_{0}$ is obtained as $H_{I}=H_{C}$. The time evolution under $H_{I}$ for the time duration $\tau$ is given by
@@
we should cancel the phonon hopping among the other modes $q_{r}\in\mathcal{S}^{c}$ as well as that between $q_{r}$ and $q_{s}$ for $q_{s}\in\mathcal{S}$ and $q_{r}\in\mathcal{S}^{c}$. For example, in order to implement the 50:50 beam-splitter between $q_{j}$ and $q_{j+1}$, we need to wait for $\tau=T_{\mathrm{50:50}}=(\pi/4)/(\kappa_{j+1,j}/2)$ while the phonon hopping involving $q_{r}\in\mathcal{S}^{c}$ should be cancelled, i.e., the phonon hopping between $q_{r_{1}}$ and $q_{r_{2}}$ for $q_{r_{1}},q_{r_{2}}\in\mathcal{S}^{c}$, that between $q_{j}$ and $q_{r}$, and that between $q_{j+1}$ and $q_{r}$ should be cancelled. Because the waiting time $\tau$ depends on the operation time of the gate applied to the modes in $\mathcal{S}$, we consider the DD to cancel the phonon hopping occurring in the course of the time propagation from $t=0$ to arbitrary time $T$. First, we explain the basic idea of the DD for a two mode case, and then extend the method to a general case with the larger number of modes.
@@ II.1.1 Two-mode DD
When only two modes are involved, the time propagation for the time duration $\tau$ corresponds to a two-mode beam-splitter $B^{\mathrm{TM}}_{1,0}(\theta_{\tau})$ given by
@@
where $I$ denotes the identity operator. It should be noted that we assume here the duration of $P_{j}$ is negligibly short. We can interpret Eq. (8) as the following sequence; (i) the phonon hopping proceeds for the first half period ($B^{\mathrm{TM}}_{1,0}(\theta_{T/2})$), (ii) $P_{j}$ is applied to flip the sign of $H_{I}$, (iii) the phonon hopping proceeds with $-H_{I}$ in the latter half, which cancels the phonon hopping in the first half as
@@ II.1.2 DD for three or more modes
The DD for the general case with more than three modes, called the concatenation, was proposed in Ref. [7]. Here, we give an algorithm to construct a pulse sequence for the concatenation. In the first step, we consider a chain of $M$ modes and group them into two sets. Although the DD works for an arbitrary grouping, we choose for the simplicity the two sets as the first half $S_{0}=\{q_{j};~{}0\leq j\leq\lfloor M/2\rfloor-1\}$ and the last half $S_{1}=\{q_{j};~{}\lfloor M/2\rfloor\leq j\leq M-1\}$. By applying $P$ to the modes in $S_{1}$ at $t=T/2$, the sign of $H_{j,k}$ for $\{j,k;~{}q_{j}\in S_{0}$,  $q_{k}\in S_{1}\}$ flips and the phonon hopping between the two sets occurring in the later half period cancels with that occurred in the first half period, i.e., $q_{j}\in S_{0}$ and $q_{k}\in S_{1}$ are decoupled. Then, in the second step, we split $S_{0}$ into $S_{00}$ and $S_{10}$ in the same manner as above and apply $P$ to the modes in $S_{10}$ at $t=T/4$ and $43T/$ 4, which flips the sign of $H_{j,k}$ for $\{j,k;~{}q_{j}\in S_{00}$,  $q_{k}\in S_{10}\}$. Similarly, we split $S_{1}$ into $S_{01}$ and $S_{11}$, and apply $P$ to the modes in $S_{11}$ at $t=T/4$ and $43T/$ 4.
@@
Finally, we need to compensate the unwanted $\pi$-phase shift. For the two-mode DD, $P_{j}$ in $U_{\mathrm{DD}}$ can be moved to the left using the relation $B^{\mathrm{TM}}_{1,0}(\theta)P_{j}=P_{j}B^{\mathrm{TM}}_{1,0}(-\theta)$ as shown by Eq. (II.1.1). For the $M$-mode DD, we can move all the $P_{j}$ in $U_{\mathrm{DD}}$ to the left in a similar manner so that we have $P_{j}^{\Pi_{j}}$, which is equal to $I$ when $\Pi_{j}$ is even and is equal to $P_{j}$ when $\Pi_{j}$ is odd. Because $\Pi_{j}$ is odd if the $q_{j}$ is in $S_{1}$, we apply $P$ to all the modes in $S_{1}$ at $t=T$. Therefore, after applying this compensation procedure, the number of $P_{j}$ is given by $\Pi_{j}=\alpha_{j}+\delta_{1,r}$, where $r$ is the right-most bit of $\alpha_{j}$ and $\delta_{1,r}$ is the Kronecker delta.
Because $P$ is not applied to $q_{0}$, which corresponds to the binary number $\alpha_{0}=00\dots 0$, we can immediately implement $B_{\mathcal{S}}$ defined by Eq. (5) using the DD described above. By using the fact that $P$ is not applied to $q_{0}$, which is associated with the binary number $\alpha_{0}=00\dots 0$, the DD described above can be immediately used to implement $B_{\mathcal{S}}$ defined by Eq. (5). First, we define $M=|\mathcal{S}^{c}|+1$ and choose a mode $q^{\prime}$ from $\mathcal{S}$. Then, by regarding $q^{\prime}$ as the zeroth mode, we apply the $M$-mode DD to the set of modes $\mathcal{S}^{\prime}=\{q^{\prime}\}\cup\mathcal{S}^{c}$. Because the phonon hoppings among the modes in $\mathcal{S}^{c}$ and that between $q^{\prime}$ and any mode in $\mathcal{S}^{c}$ are cancelled and because the choice of $q^{\prime}$ to which $P$ is not applied is arbitrary, the phonon hopping is not affected only among $\mathcal{S}$ by the DD, and consequently, $B_{\mathcal{S}}$ is securely realized.
We note that the present algorithm also works even if the roles of the subscripts ”0” and ”1” are interchanged at each step even though we assign the subscript ”1” to the subsets to which $P$ is applied to obtain the simple relation between $\alpha_{j}$ and $\Pi_{j}$. For example, at $l$-th step, we can apply $P$ to the modes in $S_{0\alpha_{l-1}}$ instead of the modes in $S_{1\alpha_{l-1}}$. As shown in III.2.1, this degree of freedom can be exploited to reduce the error in the DD.
@@
The problem on the applicability of the DD to the large $M$ can be addressed by truncating the Coulomb interaction at a certain mode separation $\eta\in\mathbb{Z}$, i.e., by omitting $H_{j,k}$ for $|j-k|>\eta$. This truncation can be rationalized because the hopping rate scales as $\kappa_{j,k}\propto(d|j-k|)^{-3}$, where $d$ denotes the distance between two neighboring ions.
For example, when $M=3\eta$, we split them into three sets $S_{0}$, $S_{1}$, and $S_{2}$, where each set contains $\eta$ modes. Because $|j-k|>\eta$ is satisfied for $\{j,k;~{}q_{j}\in S_{0},\,q_{k}\in S_{2}\}$, we do not need to consider the DD to decouple $S_{0}$ and $S_{2}$. First, we apply $P$ to the modes in $S_{1}$ at $t=T/2$, which decouples $q_{j}\in S_{0}$ and $q_{k}\in S_{2}$ form $q_{l}\in S_{1}$. Then, we split each of $S_{0}$, $S_{1}$, and $S_{2}$ into two subsets and repeat the procedure, so that $N_{\mathrm{BP}}$ is given by $2^{\beta}$ with $\beta=\mathrm{min}(\lceil\log_{2}\eta\rceil+1,\,l_{\max})$.
@@ II.1.3 Repeated-DD
As shown by Eq. (8), the two-mode DD can completely cancel the phonon hopping, but the $M$-mode DD with $M\geq 3$ cannot do so even if $T_{P}=0$. Because we associate ”1” in the subscript of $S_{\alpha_{j}}$ with the application of $P$, the unitary operator describing the 3-mode DD is given by
@@
where $B_{3}=B_{+++}$. Unlike the 2-mode DD (8), the four operators in the last line of Eq. (II.1.3) do not cancel with each other, and consequently, $U_{\mathrm{DD}}=I$ cannot be satisfied.
@@
$\hskip 14.22636pt+\mathcal{O}(\tilde{\alpha}_{\mathrm{comm}}T^{2}N_{\mathrm{BP}}^{-2}),$ (15)
where $\tilde{\alpha}_{\mathrm{comm}}=\sum_{\xi,\xi^{\prime}=\{(1,0),(2,1),(2,0)\}}(-\kappa_{\xi}\kappa_{\xi^{\prime}}/4)\|[\alpha_{\xi},\alpha_{\xi^{\prime}}]\|$, $\alpha_{\xi=(j,k)}=a_{j}^{\dagger}a_{k}+a_{j}a_{k}^{\dagger}$, and $\|A\|\geq 0$ stands for the spectral norm of $A$. Due to the relations $[a_{j},a_{k}]=0$ for all $\{j,k\}$ and $[a^{\dagger}_{j},a_{k}]=0$ for $j\neq k$, $[\alpha_{(j,k)},\alpha_{(j^{\prime},k^{\prime})}]$ becomes non-zero only when $(j=j^{\prime},k\neq k^{\prime})$ or $(j\neq j^{\prime},k=k^{\prime})$ is satisfied. Because both of $B^{\mathrm{TM}}_{j,k}(+(\bm{\theta}_{\tau})_{j,k})$ and $B^{\mathrm{TM}}_{j,k}(-(\bm{\theta}_{\tau})_{j,k})$ appear twice and cancel with each other in the last line of Eq. (II.1.3), we can obtain $U_{\mathrm{DD}}=I+\mathcal{O}(\tilde{\alpha}_{\mathrm{comm}}T^{2}N_{\mathrm{BP}}^{-2})$. Furthermore, because the coupling strength $\kappa_{j+1,j}$ becomes the largest for the nearest-neighbor modes and is given by $\kappa_{j+1,j}=\kappa_{1,0}$ for an equidistant chain of ions, $\tilde{\alpha}_{\mathrm{comm}}$ can be approximated by $(-\kappa_{1,0}^{2}/4)\sum_{\xi,\xi^{\prime}}\|[\alpha_{\xi},\alpha_{\xi^{\prime}}]\|$, where $\xi=(j+1,j)$ and $\xi^{\prime}=(j^{\prime}+1,j^{\prime})$, and consequently, the error can be rewritten as $\mathcal{O}(\kappa_{1,0}^{2}T^{2}N_{\mathrm{BP}}^{-2})$. It should be noted here that the time scale $T$ of the phonon hopping between the nearest-neighbor modes is proportional to $\kappa_{1,0}^{-1}$, e.g., the 50:50 beam-splitter between the nearest-neighbor modes can be realized with $T=(\pi/4)/(\kappa_{1,0}/2)$. Therefore, for inducing a fixed amount of phonon hopping, $\kappa_{1,0}^{2}$ cancels $T^{2}$ in the error term $\mathcal{O}(\kappa_{1,0}^{2}T^{2}N_{\mathrm{BP}}^{-2})$. Moreover, because $N_{\mathrm{BP}}$ is determined by the algorithm explained in II.1.2, there is no free parameter which we can use to reduce the error in the DD during a fixed amount of phonon hopping.
@@
which we call a repeated-DD. When $B_{3}$ is approximated by the product of $B^{\mathrm{TM}}_{j,k}((\bm{\theta}_{\tau})_{j,k}/n_{r})$ as in Eq. (II.1.3), the error, which is scaled as in $\tilde{U}_{\mathrm{DD}}^{n_{r}}=I+\mathcal{O}(\tilde{\alpha}_{\mathrm{comm}}T^{2}N_{\mathrm{BP}}^{-2}n_{r}^{-1})$, can be approximated to be $\mathcal{O}(\kappa_{1,0}^{2}T^{2}N_{\mathrm{BP}}^{-2}n_{r}^{-1})$. Therefore, by increasing $n_{r}$, we can suppress the error and obtain $\tilde{U}_{\mathrm{DD}}^{n_{r}}\simeq I$.

### v2

@@ Abstract
The local modes of trapped ions can be used to construct an analog quantum simulator and a digital quantum computer. However, the control of the phonon hopping remains difficult because it proceeds among all the local modes through the Coulomb coupling. We propose a method to cancel the phonon hopping among a given set of local modes by applying a sequence of phase shift gates implemented through the modulation of the trap potential. We analyze the error scaling in the algorithm to treat three or more modes and show that the error can be suppressed by repeating the pulse sequence. The duration of the phase shift gate in the present method can be as short as a few microseconds, which is an order of magnitude faster than the laser-based method. This short duration of the phase shift gate facilitates the suppression of the gate error. We finally show how the present method can be applied to the implementation of the beam splitter. The present method can also be applied to the simulation of bosonic systems as well as to the continuous variable encoding of quantum computing using trapped ions.
@@ I INTRODUCTION
The local-mode scheme can be used to realize the quantum simulation of, e.g., a Bose–Hubbard model [3] and a Jaynes–Cummings–Hubbard model [4], and can also be used to implement a quantum walk [6] and a boson sampling [7]. Moreover, it has been shown that a logical qubit in the CV encoding called the Gottesman–Kitaev–Preskill (GKP) state can be constructed using a local mode [8, 9, 10].
However, in the local-mode scheme, the manipulation of entanglement among two or more local modes remains difficult because the phonon hopping, a source of entanglement, originating from the Coulomb interaction, cannot be turned off. On the other hand, in the collective mode scheme [11], the modes can be entangled using a beam-splitter gate, which can be switched on only by irradiating the ions with a laser [11] or by applying the spatially and temporally varying electric potential [12].
There are several methods for the control of the phonon hopping. In the explanation below, we refer to a system having two local modes for simplicity, but the discussion can be extended to a system having more than two local modes. The first method is called the phonon blockade [13], where two local modes become off-resonant by the irradiation of one of the two local modes with laser light so that the phonon hopping can be suppressed. The method protects the phonon state of one mode, but disturbs that of the other mode being excited by the laser light.
@@
In the present study, we adopt the third method and propose to implement the $\pi$-phase shift by modulating the trap potential using an electric pulse applied to the trap electrodes and show that the fidelity of the DD can be improved. The duration of $P$ can be a few microseconds, which is shorter by an order of magnitude than that which can be achieved by the resonant sideband pulse approach. It should be noted that the present method can be applied to a system having any number of phonons. We also show that the original DD approach  [7] needs to be modified when the number of modes becomes larger than two even when the duration of $P$ is negligibly small. Finally, we show that the present method can be used to implement the beam splitter among the local modes.
@@ II.1 A sequence of phase shifts for the dynamical decoupling
We consider a chain of ions and assume that the frequencies of the oscillation along the radial direction $\tilde{x}$, perpendicular to the trap axis, are a few MHz, while the hopping rate is a few kHz. This condition allows us to adopt the local-mode picture along the $\tilde{x}$ axis. Therefore, if the number of trapped ions is $M$, we consider $M$ local modes and denote the set of those modes as $\mathcal{U}=\{q_{j};0\leq j\leq M-1\}$. The scaled coordinate of the $j$-th mode $x_{j}$ is represented using the annihilation and creation operators as $x_{j}=(a_{j}^{\dagger}+a_{j})/\sqrt{2}$. As shown in Appendix A, when the distance between adjacent ions is much larger than the displacement of the respective ions from their equilibrium positions, the Hamiltonian is composed of the on-site interaction $H_{0}$ and the Coulomb interaction $H_{C}$ as
@@
where $\omega_{0}$ in Eq. (2) is the secular frequency obtained under the influence of the Coulomb interactions. The rotating-wave approximation (RWA) has been applied to $H_{C}$ so that the fast oscillating terms such as $a_{j}a_{k}+a_{j}^{\dagger}a_{k}^{\dagger}$ can be omitted. The Hamiltonian in the interaction picture with respect to $H_{0}$ is obtained as $H_{I}=H_{C}$. The time evolution under $H_{I}$ for the time duration $\tau$ is given by
@@
we should cancel the phonon hopping among the other modes $q_{r}\in\mathcal{S}^{c}$ as well as that between $q_{r}$ and $q_{s}$ for $q_{s}\in\mathcal{S}$ and $q_{r}\in\mathcal{S}^{c}$. For example, in order to implement the 50:50 beam splitter between $q_{j}$ and $q_{j+1}$, we need to wait for $\tau=T_{\mathrm{50:50}}=(\pi/4)/(\kappa_{j+1,j}/2)$ while the phonon hopping involving $q_{r}\in\mathcal{S}^{c}$ should be canceled, i.e., the phonon hopping between $q_{r_{1}}$ and $q_{r_{2}}$ for $q_{r_{1}},q_{r_{2}}\in\mathcal{S}^{c}$, that between $q_{j}$ and $q_{r}$, and that between $q_{j+1}$ and $q_{r}$ should be canceled. Because the waiting time $\tau$ depends on the operation time of the gate applied to the modes in $\mathcal{S}$, we consider the DD to cancel the phonon hopping occurring in the course of the time propagation from $t=0$ to arbitrary time $T$. First, we explain the basic idea of the DD for a two-mode case and then extend the method to a general case with the larger number of modes.
@@ II.1.1 Two-mode DD
When only two modes are involved, the time propagation for the time duration $\tau$ corresponds to a two-mode beam splitter $B^{\mathrm{TM}}_{1,0}(\theta_{\tau})$ given by
@@
where $I$ denotes the identity operator. It should be noted that we assume here the duration of $P_{j}$ is negligibly short. We can interpret Eq. (8) as the following sequence: (i) The phonon hopping proceeds for the first half period [$B^{\mathrm{TM}}_{1,0}(\theta_{T/2})$], (ii) $P_{j}$ is applied to flip the sign of $H_{I}$, (iii) the phonon hopping proceeds with $-H_{I}$ in the latter half, which cancels the phonon hopping in the first half as
@@ II.1.2 DD for three or more modes
The DD for the general case with more than three modes, called the concatenation, was proposed in Ref. [7]. Here, we give an algorithm to construct a pulse sequence for the concatenation. In the first step, we consider a chain of $M$ modes and group them into two sets. Although the DD works for an arbitrary grouping, we choose, for simplicity, the two sets as the first half $S_{0}=\{q_{j};~{}0\leq j\leq\lfloor M/2\rfloor-1\}$ and the last half $S_{1}=\{q_{j};~{}\lfloor M/2\rfloor\leq j\leq M-1\}$. By applying $P$ to the modes in $S_{1}$ at $t=T/2$, the sign of $H_{j,k}$ for $\{j,k;~{}q_{j}\in S_{0}$,  $q_{k}\in S_{1}\}$ flips and the phonon hopping between the two sets occurring in the later half period cancels with that which occurred in the first half period, i.e., $q_{j}\in S_{0}$ and $q_{k}\in S_{1}$ are decoupled. Then, in the second step, we split $S_{0}$ into $S_{00}$ and $S_{10}$ in the same manner as above and apply $P$ to the modes in $S_{10}$ at $t=T/4$ and $43T/$ 4, which flips the sign of $H_{j,k}$ for $\{j,k;~{}q_{j}\in S_{00}$,  $q_{k}\in S_{10}\}$. Similarly, we split $S_{1}$ into $S_{01}$ and $S_{11}$, and apply $P$ to the modes in $S_{11}$ at $t=T/4$ and $43T/$ 4.
@@
Finally, we need to compensate for the unwanted $\pi$-phase shift. For the two-mode DD, $P_{j}$ in $U_{\mathrm{DD}}$ can be moved to the left using the relation $B^{\mathrm{TM}}_{1,0}(\theta)P_{j}=P_{j}B^{\mathrm{TM}}_{1,0}(-\theta)$, as shown by Eq. (II.1.1). For the $M$-mode DD, we can move all the $P_{j}$ in $U_{\mathrm{DD}}$ to the left in a similar manner so that we have $P_{j}^{\Pi_{j}}$, which is equal to $I$ when $\Pi_{j}$ is even and is equal to $P_{j}$ when $\Pi_{j}$ is odd. Because $\Pi_{j}$ is odd if the $q_{j}$ is in $S_{1}$, we apply $P$ to all the modes in $S_{1}$ at $t=T$. Therefore, after applying this compensation procedure, the number of $P_{j}$ is given by $\Pi_{j}=\alpha_{j}+\delta_{1,r}$, where $r$ is the rightmost bit of $\alpha_{j}$ and $\delta_{1,r}$ is the Kronecker delta.
Because $P$ is not applied to $q_{0}$, which corresponds to the binary number $\alpha_{0}=00\dots 0$, we can immediately implement $B_{\mathcal{S}}$ defined by Eq. (5) using the DD described above. First, we define $M=|\mathcal{S}^{c}|+1$ and choose a mode $q^{\prime}$ from $\mathcal{S}$. Then, by regarding $q^{\prime}$ as the zeroth mode, we apply the $M$-mode DD to the set of modes $\mathcal{S}^{\prime}=\{q^{\prime}\}\cup\mathcal{S}^{c}$. Because the phonon hoppings among the modes in $\mathcal{S}^{c}$ and that between $q^{\prime}$ and any mode in $\mathcal{S}^{c}$ are canceled and because the choice of $q^{\prime}$ to which $P$ is not applied is arbitrary, the phonon hopping is not affected only among $\mathcal{S}$ by the DD and, consequently, $B_{\mathcal{S}}$ is securely realized. An example for $M=3$ will be shown in Sec. III.2.3.
We note that the present algorithm also works even if the roles of the subscripts “0” and “1” are interchanged at each step, even though we assign the subscript “1” to the subsets to which $P$ is applied to obtain the simple relation between $\alpha_{j}$ and $\Pi_{j}$. For example, at the $l$-th step, we can apply $P$ to the modes in $S_{0\alpha_{l-1}}$ instead of the modes in $S_{1\alpha_{l-1}}$. As shown in Sec. III.2.1, this degree of freedom can be exploited to reduce the error in the DD.
@@
The problem of the applicability of the DD to the large $M$ can be addressed by truncating the Coulomb interaction at a certain mode separation $\eta\in\mathbb{Z}$, i.e., by omitting $H_{j,k}$ for $|j-k|>\eta$. This truncation can be rationalized because the hopping rate scales as $\kappa_{j,k}\propto(d|j-k|)^{-3}$, where $d$ denotes the distance between two neighboring ions.
For example, when $M=N\eta$ with $N\geq 2$, we divide the set of $M$ modes into $N$ subsets as $S_{0}$, $S_{1},\,\dots$, $S_{N-1}$ so that each subset contains $\eta$ modes. Because $|j-k|>\eta$ is satisfied for $\{j,k;~{}q_{j}\in S_{J},\,q_{k}\in S_{K}\}$ if $|J-K|>1$, we do not need to consider the DD to decouple $S_{J}$ from $S_{K}$ for $|J-K|>1$. Therefore, we first apply $P$ to the modes in the subsets with odd subscripts, $S_{2n+1}$, at $t=T/2$, which decouples the modes in the neighboring subsets, $q_{j}\in S_{2n}$ and $q_{k}\in S_{2(n+1)}$, from $q_{l}\in S_{2n+1}$. Then, we apply the DD to each of $S_{2n}$, $S_{2n+1}$, and $S_{2(n+1)}$, so that $N_{\mathrm{BP}}$ is given by $2^{\beta}$ with $\beta=\mathrm{min}(\lceil\log_{2}\eta\rceil+1,\,l_{\max})$.
@@ II.1.3 Repeated-DD
As shown by Eq. (8), the two-mode DD can completely cancel the phonon hopping, but the $M$-mode DD with $M\geq 3$ cannot do so even if $T_{P}=0$. Because we associate “1” in the subscript of $S_{\alpha_{j}}$ with the application of $P$, the unitary operator describing the three-mode DD is given by
@@
where $B_{3}=B_{+++}$. Unlike the two-mode DD (8), the four operators in the last line of Eq. (II.1.3) do not cancel with each other and, consequently, $U_{\mathrm{DD}}=I$ cannot be satisfied.
@@
$\hskip 14.22636pt+O(\tilde{\alpha}_{\mathrm{comm}}T^{2}N_{\mathrm{BP}}^{-2}),$ (15)
where $\tilde{\alpha}_{\mathrm{comm}}=\sum_{\xi,\xi^{\prime}=\{(1,0),(2,1),(2,0)\}}(-\kappa_{\xi}\kappa_{\xi^{\prime}}/4)\|[\alpha_{\xi},\alpha_{\xi^{\prime}}]\|$, $\alpha_{\xi=(j,k)}=a_{j}^{\dagger}a_{k}+a_{j}a_{k}^{\dagger}$, and $\|A\|\geq 0$ stands for the spectral norm of $A$. Due to the relations $[a_{j},a_{k}]=0$ for all $\{j,k\}$ and $[a^{\dagger}_{j},a_{k}]=0$ for $j\neq k$, $[\alpha_{(j,k)},\alpha_{(j^{\prime},k^{\prime})}]$ becomes nonzero only when $(j=j^{\prime},k\neq k^{\prime})$ or $(j\neq j^{\prime},k=k^{\prime})$ is satisfied. Because both $B^{\mathrm{TM}}_{j,k}(+(\bm{\theta}_{\tau})_{j,k})$ and $B^{\mathrm{TM}}_{j,k}(-(\bm{\theta}_{\tau})_{j,k})$ appear twice and cancel with each other in the last line of Eq. (II.1.3), we can obtain $U_{\mathrm{DD}}=I+O(\tilde{\alpha}_{\mathrm{comm}}T^{2}N_{\mathrm{BP}}^{-2})$. Furthermore, because the coupling strength $\kappa_{j+1,j}$ becomes the largest for the nearest-neighbor modes and is given by $\kappa_{j+1,j}=\kappa_{1,0}$ for an equidistant chain of ions, $\tilde{\alpha}_{\mathrm{comm}}$ can be approximated by $(-\kappa_{1,0}^{2}/4)\sum_{\xi,\xi^{\prime}}\|[\alpha_{\xi},\alpha_{\xi^{\prime}}]\|$, where $\xi=(j+1,j)$ and $\xi^{\prime}=(j^{\prime}+1,j^{\prime})$, and consequently, the error can be rewritten as $O(\kappa_{1,0}^{2}T^{2}N_{\mathrm{BP}}^{-2})$. It should be noted here that the time scale $T$ of the phonon hopping between the nearest-neighbor modes is proportional to $\kappa_{1,0}^{-1}$, e.g., the 50:50 beam splitter between the nearest-neighbor modes can be realized with $T=(\pi/4)/(\kappa_{1,0}/2)$. Therefore, for inducing a fixed amount of phonon hopping, $\kappa_{1,0}^{2}$ cancels $T^{2}$ in the error term $O(\kappa_{1,0}^{2}T^{2}N_{\mathrm{BP}}^{-2})$. Moreover, because $N_{\mathrm{BP}}$ is determined by the algorithm explained in Sec. II.1.2, there is no free parameter which we can use to reduce the error in the DD during a fixed amount of phonon hopping.
@@
which we call a repeated DD. When $B_{3}$ is approximated by the product of $B^{\mathrm{TM}}_{j,k}((\bm{\theta}_{\tau})_{j,k}/n_{r})$ as in Eq. (II.1.3), the error, which is scaled as in $\tilde{U}_{\mathrm{DD}}^{n_{r}}=I+O(\tilde{\alpha}_{\mathrm{comm}}T^{2}N_{\mathrm{BP}}^{-2}n_{r}^{-1})$, can be approximated to be $O(\kappa_{1,0}^{2}T^{2}N_{\mathrm{BP}}^{-2}n_{r}^{-1})$. Therefore, by increasing $n_{r}$, we can suppress the error and obtain $\tilde{U}_{\mathrm{DD}}^{n_{r}}\simeq I$.